# Equivalence and Separation between Heard-Of and Asynchronous Message-Passing Models

### v1

@@ 3 Simulating $\mathrm{HO}_{f}$ in $\mathrm{AMP}_{f}$
Finally, suppose that some protocol solves a task $T$ in $\mathrm{HO}_{f}$. Any run in $\mathrm{CFHO}_{f}$ has a corresponding run in $\mathrm{HO}_{f}$ where each process that originally crashes in some round $r$ does not crash, but instead from round $r$, (1) its messages are still lost to all other processes, and (2) it hears from all originally non-crashed processes (by Weak Liveness, there are at least nfn-f such processes). Since processes that decide in the given $\mathrm{CFHO}_{f}$ run decide based on the same information as in the constructed $\mathrm{HO}_{f}$ run, their outputs remain unchanged and moreover, non-crashed processes decide in the given run. This implies that the same protocol also solves $T$ in $\mathrm{CFHO}_{f}$, which concludes the proof of the lemma. ∎
@@ 4 From $\mathrm{AMP}_{f}$ to $\mathrm{HO}_{f}$
Figure 3: Strategy for equivalence results; $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ indicates a simulation of $\mathcal{M}_{1}$ in $\mathcal{M}_{2}$
@@ 4.1 The Silenced-Faulty Heard-Of model
The reach of process $p_{i}$ starting from round $r$, denoted $\textsc{Reach}_{i}(r,\infty)$, is $\bigcup_{s\geq r}\textsc{Reach}_{i}(r,s)$.
@@
Let $Sil$ be the set of silenced processes. Suppose that $p_{i}\in Sil$ is silenced from round $r_{i}$. Then all the processes in $\textsc{Reach}_{i}(r_{i},\infty)$ must be silenced. Indeed, if some process $p_{j}\in\textsc{Reach}_{i}(r_{i},\infty)$ is not, then by Lemma 2, $\left\lvert\textsc{Reach}_{j}(r^{\prime},\infty)\right\rvert=n$ for some $r^{\prime}>r_{i}$ and so $\left\lvert\textsc{Reach}_{i}(r_{i},\infty)\right\rvert=n$, contradicting our assumption that $p_{i}$ is silenced.
By the definition of reach, there is a round after which no message sent by a process in $\textsc{Reach}_{i}(r_{i},\infty)$ is received by a process outside the set. As we have shown for any $p_{i}\in Sil$ that $\textsc{Reach}_{i}(r_{i},\infty)\subseteq Sil$, there is a round after which no message sent by a process in $Sil$ is received by a process outside $Sil$.
@@ 4.2 From $\mathrm{AMP}_{f}$ to $\mathrm{SFHO}_{f}$
We construct a simulation system that simulates $\mathrm{AMP}_{f}$ in $\mathrm{SFHO}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$, and (2) the $i$-th simulation machine $P_{i}$ has the specification given by Algorithm 2. Roughly, to simulate sending a message in $\mathrm{AMP}_{f}$, it is echoed in every round of $\mathrm{SFHO}_{f}$ by the processes that have already received it, in the hope that all processes eventually receive it. This fails when the message’s original process is silenced in $\mathrm{SFHO}_{f}$, and the simulated run risks violating the Non-faulty Liveness condition of $\mathrm{AMP}_{f}$. So processes rely on acknowledgements to determine if their messages have reached everyone, until which they temporarily block themselves from proceeding in $\mathrm{AMP}_{f}$. Thus if they are silenced in $\mathrm{SFHO}_{f}$, they are faulty in $\mathrm{AMP}_{f}$, and the Faulty Quasi-Liveness condition allows their last message to be not received by everyone in the simulated run.
@@
2:Initialize set $old:=\emptyset$ $\triangleright$ History of messages received before $p_{i}$’s last amp-send
@@
5:Enable $\operatorname{ho-send}_{i}(seen)$ $\triangleright$ Enable send for the first round of $\mathrm{SFHO}_{f}$
6:When $\operatorname{amp-send}_{i}(m)$ occurs:
@@
12:  If $latest\neq\texttt{null}$ and $ack(j,latest)$ occurs in $M^{\prime}$ for $\geq f$ distinct values of $j\neq id_{i}$:
@@
22:  Enable $\operatorname{ho-send}_{i}(seen)$ $\triangleright$ Enable send for the next round of $\mathrm{SFHO}_{f}$
To solve a task $T$ in $\mathrm{SFHO}_{f}$, a protocol $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ can again be constructed in the same manner as in Section 3, except that now the main events of $q_{i}$ are ho-send and ho-recv, while its internal events are amp-send and amp-recv.
The simulator $q_{i}$ works as follows. The input of $q_{i}$ is also used as the input for $p_{i}$. When $p_{i}$ does an internal $\operatorname{amp-send}_{i}(m)$, $q_{i}$ broadcasts $latest=\left\langle id_{i},m,T\right\rangle$ for some time value $T$ in every subsequent round of $\mathrm{SFHO}_{f}$. Also in every round, it broadcasts all previously received messages and acknowledgements for messages of the form $\left\langle\cdot,\cdot,\cdot\right\rangle$. It waits to receive at least $f$ acknowledgements for $\left\langle id_{i},m,T\right\rangle$ before unblocking $p_{i}$ by enabling an internal $\operatorname{amp-recv}_{i}(m^{\prime})$ for every pending received message of the form $\left\langle\cdot,m^{\prime},\cdot\right\rangle$. If $p_{i}$ internally decides an output $o$, then $q_{i}$ decides $o$ as well.
Every $P_{i}$ tags every message $m$ from $p_{i}$ with the identifier $id_{i}$ and a timestamp $T$, to distinguish an $\mathrm{AMP}_{f}$ message that is sent multiple times by a process or by different processes.
@@
As a consequence, Lemma 3 implies that there are at most $f$ faulty processes in $\mathcal{P}$ in the simulated run.
In the simulated run, if a process $p_{i}$ has an $\operatorname{amp-send}(m)$ event eventually followed by an amp-recv event, it means that $q_{i}$ received at least $f$ acknowledgements to $\left\langle id_{i},m,T\right\rangle$ (for some $T$) from others before allowing further amp-recvs at $p_{i}$. Thus eventually all processes $q_{j}$ receive $\left\langle id_{i},m,T\right\rangle$ and as all non-faulty processes $p_{j}$ have infinitely many amp-recvs, $\operatorname{amp-recv}(m)$ must be one of them. Non-faulty Liveness and Faulty Quasi-Liveness hold as a consequence. Finally, it is straightforward to show that the simulated run satisfies the Integrity and No Duplicates properties.
@@
The next sections study the conditions under which a task solvable in $\mathrm{SFHO}_{f}$ is also solvable in $\mathrm{HO}_{f}$, and then apply Lemmas 1 and 4.
@@ 5 Equivalence for colorless tasks
Thus protocol $\mathcal{Q}$ solves colorless task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{HO}_{f}$, proving that for $n>2f$, colorless tasks solvable in $\mathrm{SFHO}_{f}$ are solvable in $\mathrm{HO}_{f}$. Combining this with Lemmas 1 and 4, we conclude that:
@@ 6.1 Equivalence with one fault
We now show for $f\leq 1$ and $n>2f$ that $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ solve the same set of colored tasks. When $f=0$, there are no silenced processes in $\mathrm{SFHO}_{0}$, and it is the same model as $\mathrm{HO}_{0}$ and thus $\mathrm{AMP}_{0}$ (by Lemmas 1 and 4). Thus, it suffices to consider the case $f=1$ and $n>2f=2$.
@@
1:Round $r$: Send and receive messages and change the state of $p_{i}$ accordingly. If $q_{i}$ does not have an output yet, then
2:   If $p_{i}$ decides $o$ in round $r$, then $q_{i}$ decides $o$
3:  Else
4:     If $q_{i}$ determines from its round-$r$ view that every other process has decided, then
5:       $q_{i}$ chooses an output which, combined with the others’ outputs and the input vector, satisfies the task.
@@
If $q_{i}$ is not silenced then it is guaranteed to decide using $p_{i}$. Now suppose that $q_{i}$ is silenced from round $R$ and it does not get to decide using line 2. By Lemma 3, since no other process in $\mathcal{Q}$ is silenced, all processes in $\mathcal{Q}\setminus\left\{q_{i}\right\}$ decide by some round $R^{\prime}$. Thus by Lemma 9, $q_{i}$ satisfies the condition in line 4 by round $\max(R,R^{\prime})+2$ and also decides, by Lemma 8. ∎
@@
There can be at most one process $q_{i}$ that decides using line 5. Indeed, if there were two such processes, both must have satisfied line 4, meaning each decided their outputs before the other; this is absurd. If all processes in $\mathcal{Q}$ decide their outputs in line 2, then $\mathcal{P}$ must yield output vector $O_{\mathcal{Q}}$ on run $\alpha$ in $\mathrm{SFHO}_{1}$ and thus $(I,O_{\mathcal{Q}})\in\Delta$. Otherwise, some process $q_{i}$ decides in line 5, and thus, all other processes $q_{j}$ decide in line 2. By Lemma 8, $q_{i}$’s output combined with the others’ outputs and the input vector $I$ satisfy the task and thus $(I,O_{\mathcal{Q}})\in\Delta$. ∎
@@
Combining this with Lemmas 1 and 4 gives
@@ 6.2 Separation with more than one fault
Algorithm 3 cannot be generalized to the case where $f>1$, since it is possible to have a run with two silenced processes such that they are the only ones hearing from themselves. Since the two silenced processes will never have information about each other, getting them to decide compatible outputs is a challenge due to symmetry. Indeed, this prevents $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ from coinciding for colored tasks and $f>1$. We demonstrate this with the renaming task [attiya_renaming_amp_1990].
The renaming task with initial name space of potentially unbounded size $M$ and new name space of size $N$ and $M>N$ is defined as follows. Every process initially has as input a distinct identifier from the initial name space. Every process with a $n-\bot o$utput must return a distinct output name from the new name space. To avoid trivial protocols for this task (e.g., for $N=n$, the protocol where process $p_{i}$ always chooses $i$ as output), we use the anonymity assumption on renaming protocols used in [attiya_renaming_amp_1990]. In particular, this means that all processes have the same decision function $\delta$ that maps the sequence of views at a process to an output value.
@@
For $n>2f$, the renaming task with an unbounded initial name space and a new name space of size $N=n+f$ is solvable in $\mathrm{AMP}_{f}$ [attiya_renaming_amp_1990, Theorem 7.9]. Trimming the size of the initial name space to N+n1N+n-1 clearly preserves the result. It is worth noting that the specifications of the $\mathrm{AMP}_{f}$ model in [attiya_renaming_amp_1990] differ in that all messages sent to non-faulty processes are required to be received. However it can be checked that the protocol works even if we weaken this by assuming Faulty Quasi-Liveness.
To show the separation, assume by way of contradiction that there is a protocol $\mathcal{P}$ solving the task in $\mathrm{HO}_{f}$. We show how to obtain an input vector on which $\mathcal{P}$ fails. Fix arbitrary distinct initial names $x_{1},\dots,x_{n-2}$ for processes $p_{1},\dots,p_{n-2}$. Let processes $p_{n-1}$ and $p_{n}$ have initial names $a$ and $b$ that will be chosen later. Consider a run of $\mathcal{P}$ in $\mathrm{HO}_{f}$ with input vector $(x_{1},\dots,x_{n-2},a,b)$ such that: (1) $p_{1},\dots,p_{n-2}$ hear from each other in every round, (2) $p_{n-1}$ and $p_{n}$ are not heard of by any other process in any round and (3) $p_{n-1}$ and $p_{n}$ get to hear from $p_{1},\dots,p_{n-2}$ in every round.

### v2

@@ 3 Simulating $\mathrm{HO}_{f}$ in $\mathrm{AMP}_{f}$
Finally, suppose that a protocol $\mathcal{P}$ solves a task $T$ in $\mathrm{HO}_{f}$. We show that $\mathcal{P}$ can also solve $T$ in $\mathrm{CFHO}_{f}$. Consider a run $\gamma$ of $\mathcal{P}$ in $\mathrm{CFHO}_{f}$. We can consider a run $\gamma^{\prime}$ of $\mathcal{P}$ in $\mathrm{HO}_{f}$ where every process that crashes in $\gamma$, say at the beginning of some round $r$, does not crash in $\gamma^{\prime}$ but instead from round $r$, (1) its messages are still lost to all other processes and (2) it hears from all non-crashed processes in $\gamma$ (by Weak Liveness, there are at least nfn-f such processes). By definition of $\mathrm{HO}_{f}$, all processes decide in $\gamma^{\prime}$. We show that (i) all processes that decide in $\gamma$ output the same value they decide in $\gamma^{\prime}$, meaning that the output vector of $\gamma$ is valid and (ii) all non-faulty processes in $\gamma$ decide. For (i), if process $p$ crashes in round $r$ in $\gamma$, as $\gamma$ and $\gamma^{\prime}$ are indistinguishable till round $r$ to $p$, it will decide a value in $\gamma$ if and only if it decides the same value before round $r$ in $\gamma^{\prime}$. If $p$ does not crash in $\gamma$, as $\gamma$ and $\gamma^{\prime}$ are indistinguishable, it decides in $\gamma$ the same values that it necessarily decides in $\gamma^{\prime}$. This also implies (ii). ∎
@@ 4 From $\mathrm{AMP}_{f}$ to $\mathrm{HO}_{f}$
Figure 3: Strategy for equivalence results; $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ denotes a simulation of $\mathcal{M}_{1}$ in $\mathcal{M}_{2}$
@@ 4.1 The Silenced-Faulty Heard-Of model
The reach $\textsc{Reach}_{i}(r,\infty)$ of $p_{i}$ starting from round $r$ is $\bigcup_{s\geq r}\textsc{Reach}_{i}(r,s)$.
@@
Let $Sil$ be the set of silenced processes. Suppose that $p_{i}\in Sil$ is silenced from round $r_{i}$. Then all the processes in $\textsc{Reach}_{i}(r_{i},\infty)$ must be silenced. Indeed, if some process $p_{j}\in\textsc{Reach}_{i}(r_{i},\infty)$ is not, then by Lemma 2, $\left\lvert\textsc{Reach}_{j}(r^{\prime},\infty)\right\rvert=~n$ for some $r^{\prime}>r_{i}$ and so $\left\lvert\textsc{Reach}_{i}(r_{i},\infty)\right\rvert=n$, contradicting our assumption that $p_{i}$ is silenced.
By the definition of reach, there is a round after which no message broadcast by a process in $\textsc{Reach}_{i}(r_{i},\infty)$ is received by a process outside the set. As we have shown for any $p_{i}\in Sil$ that $\textsc{Reach}_{i}(r_{i},\infty)\subseteq Sil$, there is a round after which no message broadcast by a process in $Sil$ is received by a process outside $Sil$.
@@ 4.2 From $\mathrm{AMP}_{f}$ to $\mathrm{SFHO}_{f}$
We construct a simulation system that simulates $\mathrm{AMP}_{f}$ in $\mathrm{SFHO}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$ and (2) the $i$-th simulation machine $P_{i}$ has the specification given by Algorithm 2. Roughly, to simulate broadcasting a message in $\mathrm{AMP}_{f}$, it is echoed in every round of $\mathrm{SFHO}_{f}$ by the processes that have already received it, in the hope that all processes eventually receive it. This fails when the message’s original process is silenced in $\mathrm{SFHO}_{f}$, and the simulated run risks violating the Non-faulty Liveness condition of $\mathrm{AMP}_{f}$. So processes rely on acknowledgements to determine if their messages have reached everyone, until which they temporarily block themselves from proceeding in $\mathrm{AMP}_{f}$. Thus if they are silenced in $\mathrm{SFHO}_{f}$, they are faulty in $\mathrm{AMP}_{f}$, and the Faulty Quasi-Liveness condition allows their last message to be not received by everyone in the simulated run.
@@
2:Initialize set $old:=\emptyset$ $\triangleright$ History of messages received before $p_{i}$’s last amp-bc
@@
5:Enable $\operatorname{ho-bc}_{i}(seen)$ $\triangleright$ Enable broadcast for the first round of $\mathrm{SFHO}_{f}$
6:When $\operatorname{amp-bc}_{i}(m)$ occurs:
@@
12:   If $latest\neq\texttt{null}$ and $ack(j,latest)$ occurs in $M^{\prime}$ for at least $f$ distinct values of $j\neq id_{i}$:
@@
22:  Enable $\operatorname{ho-bc}_{i}(seen)$ $\triangleright$ Enable broadcast for the next round of $\mathrm{SFHO}_{f}$
To solve a task $T$ in $\mathrm{SFHO}_{f}$, a protocol $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ can again be constructed in the same manner as in Section 3, except that now the main events of $q_{i}$ are ho-bc and ho-recv, while its internal events are amp-bc and amp-recv.
The simulator $q_{i}$ works as follows. The input of $q_{i}$ is also used as the input for $p_{i}$. When $p_{i}$ does an internal $\operatorname{amp-bc}_{i}(m)$, $q_{i}$ broadcasts $latest=\left\langle id_{i},m,T\right\rangle$ for some time value $T$ in every subsequent round of $\mathrm{SFHO}_{f}$. Also in every round, it broadcasts all previously received messages and acknowledgements for messages of the form $\left\langle\cdot,\cdot,\cdot\right\rangle$. It waits to receive at least $f$ acknowledgements for $\left\langle id_{i},m,T\right\rangle$ before unblocking $p_{i}$ by enabling an internal $\operatorname{amp-recv}_{i}(m^{\prime})$ for every pending received message of the form $\left\langle\cdot,m^{\prime},\cdot\right\rangle$. If $p_{i}$ internally decides an output $o$, then $q_{i}$ decides $o$ as well.
Every $P_{i}$ tags every message $m$ from $p_{i}$ with the identifier $id_{i}$ and a timestamp $T$, to distinguish an $\mathrm{AMP}_{f}$ message that is broadcast multiple times by a process or by different processes.
@@
Lemma 3 implies that there are at most $f$ faulty processes in $\mathcal{P}$ in the simulated run.
In the simulated run, if a process $p_{i}$ has an $\operatorname{amp-bc}(m)$ event eventually followed by an amp-recv event, it means that $q_{i}$ received at least $f$ acknowledgements to $\left\langle id_{i},m,T\right\rangle$ (for some $T$) from others before allowing further amp-recvs at $p_{i}$. Thus eventually all processes $q_{j}$ receive $\left\langle id_{i},m,T\right\rangle$ and as all non-faulty processes $p_{j}$ have infinitely many amp-recvs, $\operatorname{amp-recv}(m)$ must be one of them. Non-faulty Liveness and Faulty Quasi-Liveness hold as a consequence. Finally, it is straightforward to show that the simulated run satisfies the Integrity and No Duplicates properties.
@@
The next sections study the conditions under which a task solvable in $\mathrm{SFHO}_{f}$ is also solvable in $\mathrm{HO}_{f}$, and then apply Lemma 1 and Lemma 4.
@@ 5 Equivalence for colorless tasks
Thus protocol $\mathcal{Q}$ solves colorless task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{HO}_{f}$, proving that for $n>2f$, colorless tasks solvable in $\mathrm{SFHO}_{f}$ are solvable in $\mathrm{HO}_{f}$. Together with Lemma 1 and Lemma 4, this implies:
@@ 6.1 Equivalence with one fault
We now show for $f\leq 1$ and $n>2f$ that $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ solve the same set of colored tasks. When $f=0$, there are no silenced processes in $\mathrm{SFHO}_{0}$, and it is the same model as $\mathrm{HO}_{0}$ and thus $\mathrm{AMP}_{0}$ (by Lemma 1 and Lemma 4). Thus, it suffices to consider the case where $f=1$ and $n>2f=2$.
@@
1:Round $r$: Broadcast and receive messages and change the state of $p_{i}$ accordingly
2:If $q_{i}$ does not have an output yet, then
3:   If $p_{i}$ decides $o$ in round $r$, then $q_{i}$ decides $o$
4:   Else if $q_{i}$ determines from its round-$r$ view that every other process has decided
5:     Then $q_{i}$ chooses an output that together with the others’ outputs and the input vector, satisfies the task
@@
If $q_{i}$ is not silenced then it is guaranteed to decide using $p_{i}$. Now suppose that $q_{i}$ is silenced from round $R$ and it does not get to decide using line 3. By Lemma 3, since no other process in $\mathcal{Q}$ is silenced, all processes in $\mathcal{Q}\setminus\left\{q_{i}\right\}$ decide by some round $R^{\prime}$. Thus by Lemma 9, $q_{i}$ satisfies the condition in line 4 by round $\max(R,R^{\prime})+2$ and also decides, by Lemma 8. ∎
@@
There can be at most one process $q_{i}$ that decides using line 5. Indeed, if there were two such processes, both must have satisfied line 4, meaning each decided their outputs before the other; this is absurd. If all processes in $\mathcal{Q}$ decide their outputs in line 3, then $\mathcal{P}$ must yield output vector $O_{\mathcal{Q}}$ on run $\alpha$ in $\mathrm{SFHO}_{1}$ and thus $(I,O_{\mathcal{Q}})\in\Delta$. Otherwise, some process $q_{i}$ decides in line 5, and thus, all other processes $q_{j}$ decide in line 3. By Lemma 8, $q_{i}$’s output combined with the others’ outputs and the input vector $I$ satisfy the task and thus $(I,O_{\mathcal{Q}})\in\Delta$. ∎
@@
Combining this with Lemma 1 and Lemma 4 implies:
@@ 6.2 Separation with more than one fault
Algorithm 3 cannot be generalized to the case where $f>1$, since it is possible to have a run with two silenced processes such that they are the only ones hearing from themselves. Since the two silenced processes will never have information about each other, getting them to decide compatible outputs is a challenge due to symmetry. Indeed, this prevents $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ from coinciding for colored tasks and $f>1$. We demonstrate this with the renaming task [4].
The renaming task with initial name space of potentially unbounded size $M$ and new name space of size $N$ and $M>N$ is defined as follows. Every process initially has as input a distinct identifier from the initial name space. Every process with a $n-\bot o$utput must return a distinct output name from the new name space. To avoid trivial protocols for this task (e.g., for $N=n$, the protocol where process $p_{i}$ always chooses $i$ as output), we use the anonymity assumption on renaming protocols used in [4]. In particular, we can assume that the initial state of a process is a function of only its initial name and that all processes have the same decision function $\delta$ that maps the sequence of views at a process to an output name.
The anonymity assumption states that for any pair of processes and any input name $i$, both the processes have the same initial state when given $i$ as input. Furthermore, let $\pi$ be a permutation of $\left\{1,\dots,n\right\}$ and let $J$ and $J^{\prime}$ be two initial configurations such that for any $1\leq i\leq n$, processes $p_{i}$ in $J$ and $p_{\pi(i)}$ in $J^{\prime}$ have the same state. Let $R$ be a run of the renaming protocol starting from $J$. Consider $\pi(R)$, also a run of the protocol, starting from $J^{\prime}$ such that the sequence of events executed by $p_{i}$ in $R$ is instead executed by $p_{\pi(i)}$ in $\pi(R)$ for all $i$. Then $p_{i}$ in $R$ and $p_{\pi(i)}$ in $\pi(R)$ must have the same sequence of states.
@@
For $n>2f$, the renaming task with an unbounded initial name space and a new name space of size $N=n+f$ is solvable in $\mathrm{AMP}_{f}$ [4, Theorem 7.9]. Trimming the size of the initial name space to N+n1N+n-1 clearly preserves the result. It is worth noting that the specifications of the $\mathrm{AMP}_{f}$ model in [4] differ in that all messages sent to non-faulty processes are required to be received. However it can be checked that the protocol works even if we weaken this by assuming Faulty Quasi-Liveness.
To show the separation, assume by way of contradiction that there is a protocol $\mathcal{P}$ solving the task in $\mathrm{HO}_{f}$. We show how to obtain an input vector on which $\mathcal{P}$ fails. Fix arbitrary distinct initial names $x_{1},\dots,x_{n-2}$ for processes $p_{1},\dots,p_{n-2}$. Let processes $p_{n-1}$ and $p_{n}$ have initial names $a$ and $b$ that will be chosen later. Consider a run of $\mathcal{P}$ in $\mathrm{HO}_{f}$ with input vector $(x_{1},\dots,x_{n-2},a,b)$ such that: (1) $p_{1},\dots,p_{n-2}$ hear from each other in every round, (2) $p_{n-1}$ and $p_{n}$ are not heard of by any other process in any round, and (3) $p_{n-1}$ and $p_{n}$ get to hear from $p_{1},\dots,p_{n-2}$ in every round.